A cinema charges £8 per ticket for evening screenings and sells 250 tickets a night on average. They estimate that the price elasticity of demand for tickets is (-) 1.6.
Calculate the expected number of tickets sold if they reduce the ticket price to £7.
1
Expert's answer
2020-02-24T09:49:09-0500

Ped = % change in Qty Demanded / % change in Price

% change in price = (7-8)/8 = - 0,125 or 12.5%

If Ped = -1.6 then ticket sales will rise by 1.6 * 12.5% = 20%

20% of 250 is 50 extra tickets

Expected tickets sold rises to 300